OpenStudy (anonymous):

Select the equation of the line that passes through the point (–2, –1) and has slope 5 in point-slope form. (Points : 4) (y – 1) = 5(x – 2) (x + 2) = 5(y – 1) (x – 2) = 5(y + 1) (y + 1) = 5(x + 2)

OpenStudy (texaschic101):

point-slope form : y - y1 = m(x - x1) slope(m) = 5 (-2,-1)....x1 = -2 and y1 = -1 now we sub y - (-1) = 5(x - (-2) y + 1 = 5(x + 2)

OpenStudy (anonymous):

thank you :-)